# The Role of Social Learning and Collective Norm Formation in Fostering Cooperation in LLM Multi-Agent Systems

### v1

@@ Abstract.
A growing body of multi-agent studies with Large Language Models (LLMs) explores how norms and cooperation emerge in mixed‑motive scenarios, where pursuing individual gain can undermine the collective good. While prior work has explored these dynamics in both richly contextualized simulations and simplified game-theoretic environments, most LLM systems featuring common-pool resource (CPR) games provide agents with explicit reward functions directly tied to their actions. In contrast, human cooperation often emerges without full visibility into payoffs and population, relying instead on heuristics, communication, and punishment. We introduce a CPR simulation framework that removes explicit reward signals and embeds cultural-evolutionary mechanisms: social learning (adopting strategies and beliefs from successful peers) and norm-based punishment, grounded in Ostrom’s principles of resource governance. Agents also individually learn from the consequences of harvesting, monitoring, and punishing via environmental feedback, enabling norms to emerge endogenously. We establish the validity of our simulation by reproducing key findings from existing studies on human behavior. Building on this, we examine norm evolution across a $2\times 2$ grid of environmental and social initialisations (resource-rich vs. resource-scarce; altruistic vs. selfish) and benchmark how agentic societies comprised of different LLMs perform under these conditions. Our results reveal systematic model differences in sustaining cooperation and norm formation, positioning the framework as a rigorous testbed for studying emergent norms in mixed-motive LLM societies. Such analysis can inform the design of AI systems deployed in social and organizational contexts, where alignment with cooperative norms is critical for stability, fairness, and effective governance of AI-mediated environments.
@@ 1. Introduction
Normative reasoning and cooperation are central to decision-making in multi-agent systems (MAS), and recent advances in Large Language Models (LLMs) have enabled these themes to be studied with natural-language agents. As such systems are increasingly embedded in human contexts, they will encounter mixed-motive settings where individual incentives conflict with collective welfare. To understand cooperation in such settings, researchers have explored both complex, high-context scenarios, such as LLM agents in historical diplomacy (Hua et al., 2023; Ren et al., 2024) or virtual societies (Park et al., 2023; Warnakulasuriya et al., 2025), and simplified, game-theoretic environments that serve as testbeds for cooperative mechanisms (Piatti et al., 2024; Rivera et al., 2024; Vallinder and Hughes, 2024). While the former capture rich social dynamics, they are often governed by layered prompt designs and engineered incentives, making it difficult to isolate the mechanisms that sustain cooperation. The latter offer greater control and interpretability, yet the pathways by which LLM societies autonomously develop norms or sustain cooperation remain underexplored.
@@
The CPR game formalizes the tension between individual incentives to over-exploit a shared resource and the collective benefit of its sustainable management. The agents must manage a shared, depletable resource.
@@
Past simulation studies in CPR settings have been carefully designed to investigate cooperation dynamics in agentic societies (Piatti et al., 2024; Piedrahita et al., 2025; Backmann et al., 2025). While informative, they often diverge from real-world conditions: in human societies, individuals rarely have full visibility into their payoffs. Instead, people act based on local heuristics, and cooperation emerges over time through shared understandings, punishment, and other social mechanisms (Centola and Baronchelli, 2015). Not to mention, LLMs can learn simple strategies in their training phase to cooperate under standard models where actions are directly related to rewards. As a result, benchmarks with directly observed rewards risk eliciting behaviors that LLMs retrieve from pretraining rather than reason about, blurring the line between memorization and genuine policy formation. To bridge this gap, we introduce a framework that draws on insights from political science and institutional economics, particularly Ostrom’s institutional design principles for governing the commons (Ostrom, 1990, 2009), and from cultural evolution theory (Bowles and Gintis, 1998; Boyd and Richerson, 2002, 2009; Henrich, 2006). Our simulator makes payoffs indirect and dynamics inferential, providing a stricter test of cooperative competence under uncertainty.
@@
Figure 1 provides an overview of our framework, which comprises four modules: Harvest and Consumption, Individual Punishment, Social Learning, and Group Decision. In Harvest and Consumption, agents choose their extraction effort and daily consumption. In Individual Punishment, agents may monitor peers and punish misbehavior at a personal cost. Through Social Learning, agents adopt strategies from peers with higher payoffs (“payoff-biased social learning”), shaping their harvesting, punishment, and normative beliefs. This is the main evolutionary mechanism in our proposal, distinguishing our work from approaches where agents form opinions gradually through discourse. Finally, in the Group Decision phase, agents form collective opinions about what constitutes group-beneficial norms. Allowing agents to converse and reflect afterwards (Piatti et al., 2024) is one way to form collective opinions; however, we observed serious limitations in scaling to many agents due to the increased number of API calls. Our proposed voting mechanism for group norms is more cost-effective and scalable, requiring only two API calls per round: one to solicit opinions and another to vote on which to adopt.
@@
After carefully validating the simulation against existing human studies, we examine how group-beneficial norms evolve in agentic societies under a $2\times 2$ matrix of environmental and social initialisations: resource-rich vs. resource-scarce environments, and altruistic vs. selfish starting strategies. By comparing outcomes across different LLMs, we identify systematic differences in their tendencies toward altruism and cooperation. Moreover, we show that punishment and social learning can evolve cooperative behaviors across different LLMs. We position this framework as a testbed for probing how various models develop strategies in mixed-motive settings, and for uncovering the underlying mechanisms that sustain collective welfare.
@@
We present a CPR simulation framework in which the mapping from actions to payoffs is latent: agents observe only local, noisy outcomes and receive no clear reward signal, requiring them to infer environmental and societal dynamics (from payoff after harvest, punishment and social cues). The framework design instantiates cultural-evolutionary mechanisms, payoff-biased social learning with optional punishments, so that cooperative norms can emerge endogenously, providing a controlled testbed for comparing behavioral tendencies across LLMs in mixed-motive settings. We introduce a scalable collective-choice procedure (propose then vote) that approximates deliberation without extensive dialogue, enabling experiments with large agent populations (two API calls per agent per round). Empirically, we (i) calibrate against human CPR findings, (ii) evaluate a $2\times 2$ design (resource-rich vs. resource-scarce $\times$ altruistic vs. selfish initializations), and (iii) compare multiple LLMs, identifying systematic differences in cooperative tendencies and a mechanism that evolves group-beneficial norms across the models we study.
@@ 2.1. Norms in agentic societies
Park et al. (2023) introduced one of the first large-scale simulations of an agentic society in the Smallville sandbox environment, where LLM-driven agents navigate rich daily-life contexts. Building on this idea, subsequent work has explored normative architectures, designs for agent societies that foster the emergence of social norms to improve collective functioning. For example, Ren et al. (2024) proposed CRSEC, a four-module framework for norm emergence encompassing Creation & Representation, Spreading, Evaluation, and Compliance, while (Li et al., 2024) developed an EvolutionaryAgent that evolves cooperative norms over time. While these studies demonstrate compelling behaviours, their highly contextualised environments make it difficult to disentangle the underlying mechanisms that drive norm formation from the incidental complexity of their settings.
@@ 2.2. Norms and cooperation in repeated games
The evolution of cooperation in MAS has been extensively studied in simple two-player games. In the Donor Game, generosity can evolve via mechanisms such as reciprocity and reputation (Vallinder and Hughes, 2024), while the Stag Hunt captures the challenge of coordination on a mutually beneficial but risky choice (Liu, 2025). These games clarify foundational mechanisms but lack the complexity of multi-agent, renewable-resource dilemmas.
@@ 2.3. Common-pool resource settings
CPR games extend the social dilemma to multiple agents drawing from a rivalrous, regenerating resource. This introduces intertemporal dynamics, such as overuse leading to collapse or underuse reducing efficiency, and brings cultural-evolutionary mechanisms to the fore, including payoff-biased social learning, conformity bias, and punishment.  Piatti et al. (2024) proposed GovSim, where cooperation emerges through iterative actions, conversation, and reflection. Their “universalization” prompt improved cooperation by telling agents, e.g., “If everyone fishes more than X, the lake will be empty,” but still relied on explicit knowledge of the payoff structure.  Piedrahita et al. (2025) adapted CPR settings to study norm enforcement via sanctioning, allowing norms to adapt over time.  Backmann et al. (2025) examined CPR settings with moral imperatives in conflict with explicit incentives. In all cases, the utility function is clearly defined, such as “units harvested” or “tokens contributed to the public good”, and directly linked to actions. However, in the real world, the link between individual actions and eventual payoffs is often noisy, delayed, or hidden, so cooperation must be learned socially rather than computed from first principles.
@@ 2.4. Cultural evolution in agentic societies
Human cooperation in CPR settings is often explained through cultural-evolutionary mechanisms. Ostrom’s principles emphasise graduated sanctions, collective-choice arrangements, and monitoring over pure utility maximisation (Ostrom, 1999, 2009). Cultural evolution highlights payoff-biased learning as well as group-level selection as evolutionary mechanisms that can select for group-beneficial norms (Boyd and Richerson, 2002; Smith, 2020). Payoff-biased learning is a common learning strategy among humans. When individuals have information about the pay-offs of others, it is possible to use these cues to adaptively bias social learning, leading to evolutionary dynamics that can be very similar to natural selection (McElreath et al., 2008). When group-beneficial norms are adaptive for individuals, payoff-biased learning can create a selective force towards group-beneficial norms. Compared to literature focused on punishment (Piedrahita et al. (2025)), cultural evolution tries to answer why punishment as a costly yet group-beneficial behavior can stabilize in a population. The explanation evolutionary models provide is that the norm of sanctioning can rapidly spread locally through conformity (Henrich and Gil-White, 2001), and spread across groups through payoff-biased learning (Boyd and Richerson, 2002).
@@ 3.1.1. State, controls, and norms (per round $t$)
A single renewable stock $R(t)\!\in\![0,K]$ (carrying capacity $K$, intrinsic growth $r$) is shared by $N$ agents. Each agent $i\in\{1,\dots,N\}$ chooses an effort $e_{i}(t)\in[0,1]$, realizes a harvest $h_{i}(t)\geq 0$, consumes a fixed $c>0$, and accumulates wealth $P_{i}(t)$. For governance, agents carry a monitoring propensity $m_{i}(t)\in[0,1]$, a punishment propensity $p_{i}(t)\in[0,1]$, and an individual norm $g_{i}(t)$ (preferred cap on own harvest; for LLM agents, induced by a language prompt). The community maintains a group norm $G(t)\geq 0$, a per-agent harvest threshold that anchors enforcement. Technology and sanctions are parameterized by productivity $\alpha>0$, penalty $\beta>0$, and punisher cost $\gamma>0$. Each agent receives a private observation
@@ 3.1.2. Environment & resource dynamics
Given efforts $\{e_{i}(t)\}_{i=1}^{N}$, we assume a standard catch function, based on the effort $e_{i}(t)$ they invested in, the fishing efficiency $\alpha$, and the resources in the pool $R(t)$.
@@
so total extraction scales linearly with current stock and individual effort (Hilborn and Walters, 2013). Post-harvest stock is
@@
The logistic specification (Verhulst growth) (Bacaër, 2011) is the workhorse in renewable-resource economics and fisheries: it captures density-dependent growth with carrying capacity $K$, yields maximal surplus production at $R=K/2$, and offers a parsimonious, well-studied baseline for policy and mechanism design. We adopt it here for transparency and comparability with classic bioeconomic models.
@@ 3.1.3. Agent actions
As shown in Figure 1, the agents in our framework take four actions, as follows.
@@
Punishment and sanctioning are important for maintaining cooperation (Ostrom, 1999; Price et al., 2002; Henrich and Boyd, 2001). Based on the punitive psychological mechanism supported by empirical research, we incorporate individual punishment in the dynamics of the framework. Each agent samples a peer $j\!\neq\!i$ uniformly and inspects with probability $m_{i}(t)$. A violation occurs if $h_{j}(t)>G(t)$. Conditional on a violation, $i$ punishes $j$ with probability $p_{i}(t)$. Let $B_{i}(t)\in\{0,1\}$ be an indicator that $i$ punished someone at $t$, and $V_{i}(t)\in\{0,1\}$ that $i$ was punished. Payoff update (pre-mortality) is
@@
If $P_{i}(t\!+\!1)<0$, agent $i$ is regarded as starved and removed (thereafter $e_{i}\!=\!0$). For LLM agents, we replace rule-based punishment with in-context judgment. At decision time, the agent receives its observation $O_{i}(t)$, the current situation, and a brief summary of a few randomly sampled peers’ recent actions and outcomes. Conditioned on this textual context, the agent chooses whether—and whom—to punish, without computing a numeric violation against a threshold.
@@
We use payoff-biased social learning as a selective force on individual strategies. There is much evidence that individuals who excel tend to be imitated excessively ((Henrich and Gil-White, 2001)), which creates a selective force toward cultural strategies that yield higher payoffs (McElreath et al., 2008; Andrews et al., 2024). In this framework, agents occasionally revise their strategies and norm beliefs.
@@
Agent $i$ meets $k$ at random and adopts $s_{k}(t)$ with the pairwise-logit rule
@@
where $\bar{P}_{i}(t)$ is a payoff (e.g., an exponential moving average) and $\delta>0$ controls selection strength ( (Szabó and Fath, 2007); Eq. 71). A small mutation $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ may be added to each adopted component to maintain exploration. In this way, the high-payoff strategy and belief spread among the population. For LLM agents, social learning is not implemented via strategy copying; it is realized in-context through language about peer outcomes and the current situation.
@@
At the end of round $t$, each agent proposes a personal harvest cap $g_{i}^{\star}(t{+}1)=f_{G,i}\!\big(O_{i}(t)\big)$, yielding the proposal set $\mathcal{G}(t)=\{g_{i}^{\star}(t{+}1)\}_{i=1}^{N}$. When proposals are numeric along a single policy dimension, we update the group norm by the median-voter rule (Black, 1948): $G(t{+}1)=\operatorname{median}\!\big(\mathcal{G}(t)\big)$. In LLM implementations, we use two short prompts per agent per round: first to propose a brief natural language collective norm, then to vote over the distinct proposals. The winner is broadcast verbatim and conditions both effort selection and enforcement in round t+1t{+}1; compliance is judged in language by the agents themselves rather than by comparing actions to a numeric threshold.

### v2

@@ Abstract.
A growing body of multi-agent studies with Large Language Models (LLMs) explores how norms and cooperation emerge in mixed‑motive scenarios, where pursuing individual gain can undermine the collective good. While prior work has explored these dynamics in both richly contextualized simulations and simplified game-theoretic environments, most LLM systems featuring common-pool resource (CPR) games provide agents with explicit reward functions directly tied to their actions. In contrast, human cooperation often emerges without explicit knowledge of the payoff structure or how individual actions translate into long-run outcomes, relying instead on heuristics, communication, and enforcement. We introduce a CPR simulation framework that removes explicit reward signals and embeds cultural-evolutionary mechanisms: social learning (adopting strategies and beliefs from successful peers) and norm-based punishment, grounded in Ostrom’s principles of resource governance. Agents also individually learn from the consequences of harvesting, monitoring, and punishing via environmental feedback, enabling norms to emerge endogenously. We establish the validity of our simulation by reproducing key findings from existing studies on human behavior. Building on this, we examine norm evolution across a $2\times 2$ grid of environmental and social initialisations (resource-rich vs. resource-scarce; altruistic vs. selfish) and benchmark how agentic societies comprised of different LLMs perform under these conditions. Our results reveal systematic model differences in sustaining cooperation and norm formation, positioning the framework as a rigorous testbed for studying emergent norms in mixed-motive LLM societies. Such analysis can inform the design of AI systems deployed in social and organizational contexts, where alignment with cooperative norms is critical for stability, fairness, and effective governance of AI-mediated environments.
Key words and phrases:
@@ 1. Introduction
Normative reasoning and cooperation are central to decision-making in multi-agent systems (MAS), and recent advances in Large Language Models (LLMs) have enabled these themes to be studied with natural-language agents. As such systems are increasingly embedded in human contexts, they will encounter mixed-motive settings where individual incentives conflict with collective welfare. To understand cooperation in such settings, researchers have explored both complex, high-context scenarios, such as LLM agents in historical diplomacy hua2023waragent; ren2024emergence or virtual societies park2023generative; warnakulasuriya2025evolution, and simplified, game-theoretic environments that serve as testbeds for cooperative mechanisms piatti2024govsim; rivera2024escalation; vallinder2024cultural. While the former capture rich social dynamics, they are often governed by layered prompt designs and engineered incentives, making it difficult to isolate the mechanisms that sustain cooperation. The latter offer greater control and interpretability, yet the pathways by which LLM societies autonomously develop norms or sustain cooperation remain underexplored.
@@
Game-theory frameworks such as common pool resources games (CPR) provide a useful tool to understand the different components of cooperation in complex social-ecological systems, and help practitioners develop efficient self-governance systems ostrom1990governing. In a CPR game, the common pool resources can be accessed by a group individuals with low or no restrictions, which can lead to over-exploitation and the “tragedy of commons” hardin1968tragedy. One important goal of the game in the context of cooperation and self-governance is to establish rules, norms, or institutions under which individuals extract an appropriate amount of resources so that the common pool resources remain regenerative and that the individuals can consume the resources efficiently in the long run. The CPR game formalizes the tension between individual incentives to over-exploit a shared resource and the collective benefit of its sustainable management. The agents must manage a shared, depletable resource.
@@
Past simulation studies in CPR settings have been carefully designed to investigate cooperation dynamics in agentic societies piatti2024govsim; piedrahita2025corrupted; backmann2025ethics. While informative, they often diverge from real-world conditions: in human societies, individuals rarely have full visibility into their payoffs. Instead, people act based on local heuristics, and cooperation emerges over time through normative values, punishment, and other social mechanisms centola2015spontaneous. Not to mention, LLMs can learn simple strategies in their training phase to cooperate under standard models where actions are directly related to rewards. As a result, benchmarks with directly observed rewards risk eliciting behaviors that LLMs retrieve from pretraining rather than reason about, blurring the line between memorization and genuine policy formation. To bridge this gap, we introduce a framework that draws on insights from political science and institutional economics, particularly Ostrom’s institutional design principles for governing the commons ostrom1990governing; ostrom2009general, and from cultural evolution theory bowles1998moral; boyd2002group; boyd2009voting; henrich2006cooperation. Our simulator makes payoffs indirect and dynamics inferential, providing a stricter test of cooperative competence under uncertainty.
@@
Figure 1 provides an overview of our framework, which comprises four modules: Harvest and Consumption, Individual Punishment, Social Learning, and Group Decision. In Harvest and Consumption, agents choose their extraction effort and daily consumption. In Individual Punishment, agents may monitor peers and punish misbehavior at a personal cost. Through Social Learning, agents adopt strategies from peers with higher payoffs (payoff-biased social learning), shaping their harvesting, punishment, and normative beliefs. This is the main evolutionary mechanism in our proposal, distinguishing our work from approaches where agents form opinions gradually through discourse. Finally, in Group Decision, agents form collective opinions about what constitutes group-beneficial norms. Allowing agents to converse and reflect afterwards piatti2024govsim is one way to form collective opinions; however, we observed serious limitations in scaling to many agents due to the increased number of API calls. Our proposed voting mechanism for group norms is more cost-effective and scalable, requiring only two API calls per round: one to solicit opinions and another to vote on which to adopt. This strategy avoids multi-turn dialogue and reflection, reducing overhead relative to conversation-based norm formation.
@@
After carefully validating the framework design against existing human studies through the simulation, we examine how group-beneficial norms evolve in agentic societies under a $2\times 2$ matrix of environmental and social initialisations: resource-rich vs. resource-scarce environments, and altruistic vs. selfish starting strategies. By comparing outcomes across different LLMs, we identify systematic differences in their tendencies toward altruism and cooperation. Moreover, we show that punishment and social learning can evolve cooperative behaviors across different LLMs. We position this framework as a testbed for probing how various models develop strategies in mixed-motive settings, and for uncovering the underlying mechanisms that sustain collective welfare.
@@
We present a CPR simulation framework in which the mapping from actions to payoffs is latent, i.e., not specified to agents: they are not given an explicit reward function or payoff table, and must infer the consequences of harvesting and sanctioning from observed outcomes and social cues (e.g., from payoff after harvest, punishment and social cues). The framework design instantiates cultural-evolutionary mechanisms, payoff-biased social learning with optional punishments, so that cooperative norms can emerge endogenously, providing a controlled testbed for comparing behavioral tendencies across LLMs in mixed-motive settings. We introduce a scalable collective-choice procedure (propose then vote) that approximates deliberation without extensive dialogue, enabling experiments with large agent populations (two API calls per agent per round).
@@ 2.1. Norms in agentic societies
park2023generative introduced one of the first large-scale simulations of an agentic society in the Smallville sandbox environment, where LLM-driven agents navigate rich daily-life contexts. Building on this idea, subsequent work has explored normative architectures, designs for agent societies that foster the emergence of social norms to improve collective functioning. For example, ren2024emergence proposed CRSEC, a four-module framework for norm emergence encompassing Creation & Representation, Spreading, Evaluation, and Compliance, while li2024agent developed an EvolutionaryAgent that evolves cooperative norms over time. While these studies demonstrate compelling behaviours, their highly contextualised environments make it difficult to disentangle the underlying mechanisms that drive norm formation from the incidental complexity of their settings.
@@ 2.2. Norms and cooperation in repeated games
[…]
@@ 2.3. Common-pool resource settings
[…]
@@ 2.4. Cultural evolution in agentic societies
Human cooperation in CPR settings is often explained through cultural-evolutionary mechanisms. Ostrom’s principles emphasise graduated sanctions, collective-choice arrangements, and monitoring over pure utility maximisation ostrom1999design; ostrom2009general. Cultural evolution highlights payoff-biased learning as well as group-level selection as evolutionary mechanisms that can select for group-beneficial norms boyd2002group; smith2020cultural. Payoff-biased learning is a common learning strategy among humans. When individuals have information about the pay-offs of others, it is possible to use these cues to adaptively bias social learning, leading to evolutionary dynamics that can be very similar to natural selection mcelreath2008beyond. When group-beneficial norms are adaptive for individuals, payoff-biased learning can create a selective force towards group-beneficial norms. Compared to literature focused on punishment piedrahita2025corrupted, cultural evolution asks why costly sanctioning behavior can stabilize in a population. One explanation is that sanctioning practices can spread locally through conformity henrich2001evolution, and spread across groups through payoff-biased learning boyd2002group.
@@ 3.1.1. State, controls, and norms (per round $t$)
A single renewable stock $R(t)\!\in\![0,K]$ (carrying capacity $K$, intrinsic growth $r$) is shared by $N$ agents. Each agent $i\in\{1,\dots,N\}$ chooses an effort $e_{i}(t)\in[0,1]$, realizes a harvest $h_{i}(t)\geq 0$, consumes a fixed $c>0$, and accumulates wealth $P_{i}(t)$. For governance, agents carry a monitoring propensity $m_{i}(t)\in[0,1]$, a punishment propensity $p_{i}(t)\in[0,1]$, and a personal normative belief $g_{i}(t)$ (preferred cap on own harvest; for LLM agents, induced by a language prompt). Here, personal normative belief is introduced to denote an agent’s internalized view of appropriate behavior (what it thinks should be done). The community maintains a group norm $G(t)\geq 0$, a per-agent harvest threshold that anchors enforcement. In an abstract sense, this represents the shared, collectively adopted expectation that anchors coordination and enforcement. Technology and sanctions are parameterized by productivity $\alpha>0$, penalty $\beta>0$, and punisher cost $\gamma>0$. Each agent receives a private observation
@@ 3.1.2. Environment & resource dynamics
Given efforts $\{e_{i}(t)\}_{i=1}^{N}$, we assume a standard catch function based on the effort $e_{i}(t)$ they invested, the fishing efficiency $\alpha$, and the resources in the pool $R(t)$.
@@
so total extraction scales linearly with current stock and individual effort hilborn2013quantitative. Post-harvest stock is
@@
The logistic specification (Verhulst growth) bacaer2011verhulst is the workhorse in renewable-resource economics and fisheries: it captures density-dependent growth with carrying capacity $K$, yields maximal surplus production at $R=K/2$, and offers a parsimonious, well-studied baseline for policy and mechanism design. We adopt it here for transparency and comparability with classic bioeconomic models.
@@ 3.1.3. Agent actions
As shown in Fig. 1, the agents in our framework take four actions, as follows.
@@
Punishment and sanctioning are important for maintaining cooperation ostrom1999design; price2002punitive; henrich2001people. Based on the punitive psychological mechanism supported by empirical research, we incorporate individual punishment in the dynamics of the framework. Each agent samples a peer $j\!\neq\!i$ uniformly and inspects with probability $m_{i}(t)$. A violation occurs if $h_{j}(t)>G(t)$. Conditional on a violation, $i$ punishes $j$ with probability $p_{i}(t)$. Let $B_{i}(t)\in\{0,1\}$ be an indicator that $i$ punished someone at $t$, and $V_{i}(t)\in\{0,1\}$ that $i$ was punished. Payoff update (pre-mortality) is
@@
If $P_{i}(t\!+\!1)<0$, agent $i$ is regarded as starved and removed (thereafter $e_{i}\!=\!0$). For LLM agents, we replace rule-based punishment with in-context judgment. At decision time, the agent receives its observation $O_{i}(t)$, the current situation, and a brief summary of a few randomly sampled peers’ recent actions and outcomes. Conditioned on this, the agent chooses whether—and whom—to punish, without computing a numeric violation against a threshold.
@@
We use payoff-biased social learning as a selective force on individual strategies. There is much evidence that individuals who excel tend to be imitated excessively (henrich2001evolution), which creates a selective force toward cultural strategies that yield higher payoffs mcelreath2008beyond; andrews2024cultural. In this framework, agents occasionally revise their strategies and norm beliefs.
@@
Agent $i$ meets $k$ at random and adopts $s_{k}(t)$ with the logit rule
@@
where $\bar{P}_{i}(t)$ is a payoff (e.g., an exponential moving average) and $\delta>0$ controls selection strength ( szabo2007evolutionary; Eq. 71). A small mutation $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ may be added to each adopted component to maintain exploration. In this way, the high-payoff strategy and belief spread among the population. For LLM agents, social learning is not implemented via strategy copying; it is realized in-context through language about peer outcomes and the current situation.
@@
At the end of round $t$, each agent proposes a personal harvest cap $g_{i}^{\star}(t{+}1)=f_{G,i}\!\big(O_{i}(t)\big)$, yielding the proposal set $\mathcal{G}(t)=\{g_{i}^{\star}(t{+}1)\}_{i=1}^{N}$. When proposals are numeric along a single policy dimension, we update the group norm by the median-voter rule (black1948rationale): $G(t{+}1)=\operatorname{median}\!\big(\mathcal{G}(t)\big)$. In LLM implementations, we use two short prompts per agent per round: first to propose a brief natural language collective norm, then to vote over the distinct proposals. The winner is broadcast verbatim and conditions both effort selection and enforcement in round t+1t{+}1; compliance is judged in language by the agents themselves rather than by comparing actions to a numeric threshold. By contrast, dialogue-based norm formation typically requires additional communication and reflection turns per round, increasing API overhead and limiting horizon and population size.